OpenStudy (anonymous):

Find the sum: 15 + (-27) answer?

OpenStudy (saifoo.khan):

15 -27 -12

OpenStudy (anonymous):

-12

jimthompson5910 (jim_thompson5910):

15 + (-27) is the same as saying 15 - 27 So subtract 27 from 15 to get -12 Note: 27 - 15 = 12, so 15 - 27 = -12 So 15 + (-27) = -12

OpenStudy (anonymous):

THANK U!